OpenStudy (anonymous):

The work of a student to solve the equation 2(3x − 4) = 8 + 2x + 4 is shown below: Step 1: 2(3x − 4) = 8 + 2x + 4 Step 2: 5x − 6 = 12 + 2x Step 3: 5x − 2x = 12 + 6 Step 4: 3x = 18 Step 5: x = 6 In which step did the student first make an error and what is the correct step? Step 2; 6x − 6 = 2(6 + x + 2) Step 2; 6x − 8 = 12 + 2x Step 3; 5x − 2x = 12 − 6 Step 3; 5x + 2x = 12 + 6

OpenStudy (haseeb96):

Step 2 ; 6x-8 =12 +2x

OpenStudy (anonymous):

Step 2; 6x − 8 = 12 + 2x

OpenStudy (here_to_help15):

lol today laziness hit me sorry ;)